OpenStudy (anonymous):

Can anyone help me solve this? 2/3(6x-18)+3/4(8x+32)=-18

OpenStudy (anonymous):

Is this your equation? \[\frac{ 2 }{ 3\left( 6x-18 \right) }+\frac{ 3 }{ 4\left( 8x+32 \right) }=-18\] or \[\frac{ 2 }{ 3 }\left( 6x-18 \right)+\frac{ 3 }{ 4 }\left( 8x+32 \right)=-18\] please clarify your problem. :)